1.4 Consider the following financial information for a company (in millions of dollars)

Partial income statement (latest year ending December 31, 2017):

Sales                                                                $900

Cost of goods sold                                          ($350)

Other expenses                                              ($200)

Depreciation                                                    ($120)

Interest expenses                                           ($30)

Partial balance sheet (as of December 31)

                                                              2016                             2017

Cash                                                    $30                                  $20

Receivables                                         $200                                $250

Inventories                                           $220                                $210

Net fixed assets                                   $500                                $520

Accounts payable                               $100                                 $120

Accruals                                                $80                                 $100

Short-term debt                                     $50                                  $10

Long-term debt                                      $270                                $270

Net income after taxes in 2017 = $150

Stockholders' equity in 2016=$450             Stockholders' equity in 2017=$500

dividend per share = dividend paid in 2017/no. of shares outstanding = $100/10 = $10

fixed assets purchased during 2017 = $520 + $120 - $500 = $140

e) What was the working capital in 2016 and 2017 (assume all cash was used for operations)?

f) Set up the company's cash flow statement for the latest year. What was its net cash flow? What was its operating cash flow, free cash flow and free cash flow to equity?

1.5 Using the information from Problem 1.4, calculate the following financial ratios for 2017:

a) Current ratio and quick ratio;

b) Inventory turnover ratio, days sales outstanding, fixed assets turnover ratio, total assets turnover ratio (Hint: Use beginning and end-of-year averages, whenever ratios combine balance sheet and income statement figures);

c) Debt (leverage) ratio, debt to equity ratio, interest coverage ratio;

d) Net profit margin, return on assets, return on equity, dividend payout ratio;

e) Assuming the company's common stock is trading at $ 150 per share, calculate its price/earnings ratio, market/book ratio, earnings yield and dividend yield.
Answer #1

Working Capital = Current Assets - Current Liabilities

Working Capital for 2016 = Current Assets (2016) - Current Liabilities (2016)
= 30 + 200 + 220 - 100 - 80 - 50
= 450 - 230
Working Capital for 2016 = $220 million

Working Capital for 2017 = Current Assets (2017) - Current Liabilities (2017)
= 20 + 250 + 210 - 120 - 100 - 10
= 480 - 230
Working Capital for 2017 = $250 million

Particulars Amount (in million)
Operating Activities
Net Income 150
Add: Depreciation 120
Changes in Working Capital
(Increase)/Decrease in Receivables (50)
(Increase)/Decrease in Inventories 10
Increase/(Decrease) in Accounts Payable 20
Increase/(Decrease) in Accruals 20
Total Change in Working Capital 0
Cash Flow from Operating Activities 270
Investing Activities
Purchase of Fixed Assets (140)
Cash Flow from Investing Activities (140)
Financing Activities
Increase/(Decrease) in Debt (40)
Dividend Paid (100)
Cash Flow from Financing Activities (140)
Net Cash (10)
Cash at the Beginning of the Year 30
Cash at the Ending of the Year 20
Particulars Amount (in million)
Operating Activities
Net Income 150
Add: Depreciation 120
Changes in Working Capital
(Increase)/Decrease in Receivables (50)
(Increase)/Decrease in Inventories 10
Increase/(Decrease) in Accounts Payable 20
Increase/(Decrease) in Accruals 20
Total Change in Working Capital 0
Cash Flow from Operating Activities 270
Investing Activities
Purchase of Fixed Assets (140)
Cash Flow from Investing Activities (140)
Financing Activities
Increase/(Decrease) in Debt (40)
Dividend Paid (100)
Cash Flow from Financing Activities (140)
Net Cash (10)
Cash at the Beginning of the Year 30
Cash at the Ending of the Year 20

Net Cash Flow is ($10 million)
Cash Flow from Operating activities is $270 million
Cash Flow from Investing activities is ($140 million)
Cash Flow from Financing activities is ($140 million)

Financial Ratios for 2017

a) Current Ratio = Current Assets/Current Liabilities
=  480/230
Current Ratio = 2.09

Quick Ratio = Current Assets - Inventories/ Current Liabilities
= 480-210/230
= 270/230
Quick Ratio = 1.17

b) Inventory Turnover Ratio = Cost of Goods Sold / Average Inventory
= 350/215
Inventory Turnover Ratio = 1.63

Days Sales Outstanding = (Average Receivables/Sales) * 365
= (225/900)*365
= 0.25*365
Days Sales Outstanding = 91 days

Fixed Asset Turnover Ratio = Sales/Average Fixed Assets
= 900/510
Fixed Asset Turnover Ratio = 1.76

Total Asset Turnover Ratio = Sales/Average Total Assets
= 900/975
Total Asset Turnover Ratio = 0.92

c) Debt Ratio = Total Debt/Total Assets
= (270+10)/1000
= 280/1000
Debt Ratio = 0.28

Debt/Equity Ratio = Total Debt/Total Equity
= (270+10)/500
= 280/500
Debt/Equity Ratio = 0.56

Interest Coverage Ratio = EBIT/Interest Expense
= 230/30
Interest Coverage Ratio = 7.67

d) Net Profit Margin = Net Profit After Tax/Sales
= 150/900
  Net Profit Margin = 16.67%

Return on Assets = Net Profit After Tax/Average Total Assets
= 150/975
Return on Assets = 15.38%

Return on Equity = Net Profit After Tax/Average Equity
= 150/475
Return on Equity = 31.58%

Dividend Payout Ratio = Dividend Paid/Net Profit After Tax
= 100/150
Dividend Payout Ratio = 66.67%

e) Price/Earnings Ratio = Market Price per share/Earnings Per Share
= 150/15 (EPS = 150/10 = 15)
Price/Earnings Ratio = 10

Market/Book Ratio = Market Price per share/Book Value Per Share
= 150/50 (BV = 500/10 = 50)
Market/Book Ratio = 3

Earnings Yield = 1/PE Ratio
= 1/10
Earnings Yield = 10%

Dividend Yield = Dividend Per Share/Market Price per share
= 10/150
Dividend Yield = 6.67%